A quart of parkers milk contains a mean of 39 grams of butterfat with a standard deviation of 2 grams. If the butterfat is normally distrubuted find the probability that the quart of this brand of milk chosen at random will contain A. 39-43 grams of butterfat of B. 36-39 grams of butterfat
Answer by stanbon(75887)

A quart of parkers milk contains a mean of 39 grams of butterfat with a standard deviation of 2 grams. If the butterfat is normally distrubuted find the probability that the quart of this brand of milk chosen at random will contain
A. 39-43 grams of butterfat
z(39) = (39-39)/2 = 0
z(43) = (43-39)/2 = 2
----
P(39<= x <=43) = P(0<= x <=2) = 0.4772
--------------------------------------------
B. 36-39 grams of butterfat
z(36) = (36-39)/2 = -3/2
z(39) = 0
---
P(36<= x <=39) = P(-3/2<= z <=0) = 0.4332
============================================
Cheers,
Stan H.
============